Charlie's sister is 3 less than 3 times c if charlie's sister is 9 years old,how old is c
Answer by stanbon(75887)

Charlie's sister is 3 less than 3 times c if charlie's sister is 9 years old,how old is c
Equations:
s = 3*c-3
s = 9 yrs
-----------------------
Solve for "c":
c = (s+3)/3
c = (12)/3 = 4 yrs
---------------------
Cheers,
Stan H.
----------